OpenStudy (anonymous):

solve for x where x is a real number (x+1)^2(x-2) + (x+1)(x-2)^2 =0

OpenStudy (anonymous):

Solve for x over the real numbers: (-2+x)^2 (1+x)+(-2+x) (1+x)^2 = 0 Write the cubic polynomial on the left hand side in standard form. Expand out terms of the left hand side: 2-3 x-3 x^2+2 x^3 = 0 Factor the left hand side. The left hand side factors into a product with three terms: (-2+x) (1+x) (-1+2 x) = 0 Solve each term in the product separately. Split into three equations: -2+x = 0 or 1+x = 0 or -1+2 x = 0 Look at the first equation: Solve for x. Add 2 to both sides: x = 2 or 1+x = 0 or -1+2 x = 0 Look at the second equation: Solve for x. Subtract 1 from both sides: x = 2 or x = -1 or -1+2 x = 0 Look at the third equation: Isolate terms with x to the left hand side. Add 1 to both sides: x = 2 or x = -1 or 2 x = 1 Solve for x. Divide both sides by 2: Answer: | | x = 2 or x = -1 or x = 1/2

OpenStudy (anonymous):

\[(x+1)^2(x-2)+(x+1)(x-2)^2 = 0 \]

OpenStudy (anonymous):

=)

OpenStudy (anonymous):

thank you!